OpenStudy (anonymous):

Last year, Elsa deposited $7000 into an account that paid 3% interest per year and $9000 into an account that paid 11% interest per year. No withdrawals were made from the accounts.

OpenStudy (anonymous):

What was the percent interest for the total deposited?

OpenStudy (anonymous):

what was the total interest erned at the end of 1 year

OpenStudy (yanjiayu):

(7000*3%+9000*11%)/(7000+9000)=7.5% the interest earned at the end of 1 year is 7000*3%+9000*11%=1200$

OpenStudy (anonymous):

thank you so much . can u help me with another one

OpenStudy (anonymous):

Suppose 11 quarts of pure antifreeze is mixed with 5 quarts of a solution that is 16% antifreeze. Answer the questions below. Do not do any rounding.

OpenStudy (anonymous):

how many quarts of antifreeze are in the resulting mixture What percentage of the resulting mixture is antifreeze? %

OpenStudy (yanjiayu):

antifreeze is 11+5*16%=11.8 quarts 11.8/(11+5)=73.75%

OpenStudy (anonymous):

thank you so much . man i really appreciate the help

OpenStudy (yanjiayu):

You are welcome.